### v1

# Developing a Reliable, General-Purpose Hallucination Detection and Mitigation Service: Insights and Lessons Learned

Abstract
Hallucination, a phenomenon where large language models (LLMs) produce output that is factually incorrect or unrelated to the input, is a major challenge for LLM applications that require accuracy and dependability. In this paper, we introduce a reliable and high-speed production system aimed at detecting and rectifying the hallucination issue within LLMs. Our system encompasses named entity recognition (NER), natural language inference (NLI), span-based detection (SBD), and an intricate decision tree-based process to reliably detect a wide range of hallucinations in LLM responses. Furthermore, our team has crafted a rewriting mechanism that maintains an optimal mix of precision, response time, and cost-effectiveness. We detail the core elements of our framework and underscore the paramount challenges tied to response time, availability, and performance metrics, which are crucial for real-world deployment of these technologies. Our extensive evaluation, utilizing offline data and live production traffic, confirms the efficacy of our proposed framework and service.
Developing a Reliable, General-Purpose Hallucination Detection and Mitigation Service: Insights and Lessons Learned
Song Wang, Xun Wang, Jie Mei, Yujia Xie, Sean Muarray, Zhang Li, Lingfeng Wu, Si-Qing Chen , Wayne Xiong Microsoft {sonwang, xunwang, jimei, yujiaxie,}@microsoft.com {murraysean, z67li, lingfw, sqchen, weixi}@microsoft.com
1 Introduction
In the rapidly evolving landscape of natural language processing (NLP), large language models (LLMs) have marked a significant leap forward, unlocking new horizons of capabilities and potentials. However, alongside their remarkable advancements, LLMs bring forth substantial challenges, with "hallucination" standing out as a particularly problematic issue. Hallucination in this context refers to instances when an LLM produces output that is either factually incorrect or not anchored in the supplied input, thus compromising the model’s reliability and the credibility of its applications.
The importance of confronting and mitigating hallucinations in LLM deployments cannot be overstated. It is foundational not only to preserving the integrity of outputs across various applications but also to averting the repercussions of misinformation in sensitive domains. Against this backdrop, our work delves into the nuances of hallucinations within LLMs, placing a special emphasis on intrinsic hallucinations, i.e. errors that can be checked against reference inputs due to their presence in more controlled settings.
A major challenge in hallucination detection and mitigation stems from the complexity and diversity of the phenomenon Ji et al. (2023); Huang et al. (2023); Tonmoy et al. (2024). Hallucinations can vary in their forms and causes depending on the context, model, task, and domain. Therefore, we conducted a comprehensive study to understand the problems by reviewing public benchmark datasets, and also analyzing real hallucinated examples reported by customers and observed during human evaluations. With that, we developed a comprehensive hallucination taxonomy and human labeling guidelines. Furthermore, we developed a GPT4-based detection method, which turns to be very competitive when compared with human annotators.
We aim to develop a production system that provides a general-purpose, fast and accurate hallucination detection and mitigation service with the existing state-of-the-art technologies. To this end, we propose our federated hallucination detection and mitigation framework, as shown in Figure 1, which includes multi-source detection, iterative rewriting and multi-source verification. The multi-source detection module is designed to be diverse and inclusive. It ensembles multiple sources, utilizing named entity recognition (NER), natural languge inference (NLI), span-based-detection (SBD) and other techniques, to recall as much hallucination errors as possible. Based on the AI feedback, we will construct a rewriting prompt dynamically and ask GPT4 to fix the errors with minimal changes. This process can be repeated until there is no hallucination errors or it reaches to a pre-defined rewriting count threshold. Then a multi-source verification module which is similar to the detection module but is optimized for precision, will be applied to further block any remaining hallucinated contents.
After overcoming various challenges, the proposed system has achieved a great balance of accuracy, latency and cost in the production and is being integrated into the customers’ LLM-based applications to reduce and mitigate the escalations about the hallucination issues. Furthermore, we have analyzed the service’s performance on offline benchmark dataset and on the real production traffic, which further demonstrate its effectiveness in serving real-world LLM applications.
In this paper, we detail the key components in our service and highlight critical challenges, lessons and potential limitations, which we believe will benefit future research in this field.
2 Related Works
Hallucination Taxonomy
A widely-adopted classification of hallucination is the intrinsic-extrinsic dichotomyDziri et al. (2021); Huang et al. (2021). Intrinsic hallucination occurs when LLM outputs contradict the provided input, such as prompts. Conversely, extrinsic hallucination occurs when LLM outputs cannot be verified by the information in the input. Recently, reasearchers have proposed more fine-grained taxonomies Pagnoni et al. (2021); Mishra et al. (2024).
We largely followed the categories in Pagnoni et al. (2021) with modifications that reflect the nature and causes of hallucinations in LLM outputs and we also developed a guidelines based on this taxonomy for annotators.
Hallucination Detection
Conventional methods of detecting hallucination can be classified into two types: token-based and sentence-based. The former aims to find hallucinated tokens while the latter is to identify the sentences with hallucinations. Various methods have been developed for identifying hallucinations and most of them leverage the pre-trained LLMs fine tuned on task-specific data Liu et al. (2021); Dziri et al. (2021); Cao et al. (2021); Zha et al. (2023). More recently, LLM with prompt-based methods are also widely used (Manakul et al., 2023; Lei et al., 2023). Though the prompt-based methods require little to no tuning and have competitive performance, they tend to have higher cost and high latency.
In our study, we’ve harnessed the strengths of both methods by employing LLM-based detection for labeling data and creating an ensemble of tailored traditional models such as NER, NLI, and span-based sequence labeling. Our designed detection service conducts a detailed examination of the input text’s semantic and syntactic attributes, allowing it to detect various kinds of hallucinations across different granularities and categories.
Hallucination Mitigation
Hallucination mitigation is to correct the identified hallucinations in the generated responses by LLMs. There are many other ways to reduce hallucination during post-generation. For example RARR Thorne and Vlachos (2021) trained a T5 model using retrieved evidence to generate corrected responses. More recently, researchers have been leveraging LLMs to better utilize hallucination feedback and generate corrections Mündler et al. (2024); Dhuliawala et al. (2023); Lei et al. (2023).
In our work, our rewriter is also LLM-based, leveraging LLM’s self refinement through feedback and reasoning. The key difference is that we have to take into consideration the cost and latency, which demands fewer output tokens, while ensuring the mitigation performance.
Table 1: Hallucination taxonomy and examples. We largely followed the categories in Pagnoni et al. (2021) with modifications that reflect the nature and causes of hallucinations in LLM output generation.
Table 2: Hallucination distribution
3 Hallucination Taxonomy and GPT4-based detection
3.1 Hallucination Taxonomy
It’s important to have a comprehensive understanding of hallucinations and we have dedicated significant effort to developing a hallucination taxonomy by manually analyzing various hallucinated documents.
In Table 1, we listed our taxonomy and examples of hallucinations, and in Table 2, we listed the distribution of different kinds of hallucinations based on human annotations over 34 document summaries. As is shown in the table, hallucinations are of different types and root causes. A majority of them are due to semantic frame errors but discourse errors and pragmatic errors also exist. The variety of hallucinations make the detection a huge challenge. Based on this taxonomy, we define our detailed hallucination guidelines, which improves the inter-annotator agreement in our study.
As stated, we mainly focus on intrinsic hallucinations, which are the errors that can be verified from the source document. While most of the existing hallucination detection solutions are entity-based which treat new entities in the generated summary as hallucinations, from table 2, we found only a small portion,  5% of hallucinations are attributed to new entities. This observation motivates us to develop a federated solution that extends NER based methods to recall more hallucination errors.
3.2 GPT4-based Evaluation
We developed a GPT4-based hallucination detection where we first split the LLM responses into individual sentences and instruct GPT4 to check its factuality sentence by sentence against the source document. This method outperforms on existing methods on multiple datasets (e.g. results on SummAC in Table 4). Other works Lei et al. (2023); Wei et al. (2024) have also developed GPT4-based detection by leveraging LLM’s reasoning capabilities. This GPT4-based detection method plays multiple important roles in our project such as labeling data for training NLI and SBD models and being used in online monitoring of hallucination pipelines in production. However, this method is not suitable to be used in our production system due to its high latency and cost.
3.3 GPT4-based Detection vs Human Annotations
Hallucination annotation is both laborious and costly, requiring clear guidance and extensive background knowledge. We conducted comprehensive studies to understand the gap between human annotation and our designed GPT4 prompt. In table 3, we list our evaluation results based on 20 hallucinated summaries, where GPT4 and humans disagree. We found GPT4 tends to produce higher false positive rates while humans tend to produce high false negative rate, and overall, GPT4 is more accurate than human.
False Positive False Negative Total
Annotators 1 10 11
GPT4 4 5 9
Table 3: Comparison of GPT4 and human
4 Federated hallucination detection and mitigation
In this section, we explain how to develop the components in the proposed federated hallucination detection and mitigation framework as shown in Figure 1.
Figure 1: Federated hallucination detection and mitigation system
4.1 Hallucination Detection
NER-based detection
Named Entity Recognition (NER) aims to identify and categorise key information (entities) in text. By applying NER analysis to the input data, we can spot possible entities that are present in the LLM outouts but not supported by the source document, i.e. hallucinations. We are using an well-known NER service [^1] Azure AI Service: https://azure.microsoft.com/en-us/products/ai-services/ai-language which will return both entity types and the confidences. Figure 2 shows the hallucination distribution over entity types in the generated LLM outputs in a human analysis. Similarly, we conducted the NER analysis on benchmark datasets in target domain, to decide the entity types and confidence thresholds in our NER detection implementation.
Figure 2: Hallucination distribution over entity types
NLI-based detection
Given two input text snippets, a premise and a hypothesis, the task of natural language inference (NLI) aims to predict the relationships: entailment, contradiction or neural. In principle, this is exactly what hallucination detection aims at. However, the premises and hypotheses in most existing NLI datasets (Bowman et al., 2015; Williams et al., 2018; Nie et al., 2020; Schuster et al., 2021) are short (one or two sentences). Therefore, we include a new document-sentence dataset (Kamoi et al., 2023) and use GPT4 to label a new dataset consisting of a diverse set of document-summary pairs. Finally, we finetune the pre-trained encoder of DeBERTa He et al. (2021) with the above datasets, and this model detects hallucinations based on semantic relationship between the document and summary.
Span-based detection
NLI is a sentence level hallucination detection while NER is restricted to a predefined set of named entities, and we further explore more fine-grained hallucination detection. Specifically, we train a token level hallucination detection model to provide more fine-grained AI feedback, e.g. hallucinated text spans. We start with the dataset labeled for NLI model and further ask GPT4 to highlight the hallucinated text spans when it is hallucinated. As shown in Figure 3, we start with a pre-trained replaced token detection head from DeBERTa He et al. (2021), and adapt it with GPT-4 generated data like a sequence labeling task. We refer to this model as Span-based detection model (SBD).
Figure 3: Span based hallucination detection
Multi-source ensemble
We believe the combination of multi-sources is necessary to further boost performance with minimal cost. For this part, we collected about 10k training examples from the target traffic, where the label (hallucinated vs not-hallucinated) is produced by GPT4 detection and the features are outputs from NER, NLI and SBD detection methods. We adopt the Gradient boosting decision tree to leverage the diverse AI feedback and finetune a model using scikit-learn to generate a single numerical value indicating whether the response or summary is hallucinated.
4.2 GPT4-based rewriting
With AI feedback from different sources, we compose a prompt based on a pre-defined template and ask GPT4 to fix it. Initially, we devise prompt version v1 that does an exhaustive chain-of-thoughts reasoning and does a full rewriting, which gives better performance but higher latency. Later on, to meet the latency requirement in the product, we developed rewriting prompt v2, which reduces chain-of-thought reasoning and then skip the rewriting if not hallucinated or otherwise only rewrite the hallucinated sentences instead of the full response.
5 Main Results
5.1 Detection Results
We compare the different detection methods’ performance on both internal and public benchmark datset. As shown in Table 4, NER tends to have lower recall but relatively higher precision in public SummAC Laban et al. (2022). It’s notable that of the three tailored models, the SBD method outperforms the rest in all metrics, showcasing the effectiveness of detailed detection - an insight that is valuable to highlight.
Table 4: Performance of different methods on the internal benchmark and public SummAC.
5.2 Mitigation Results
We define mitigation rate as percentage of hallucination being fixed and use GPT4 output tokens length as proxy to the latency and cost. As shown in Table 5, prompt v2 is more token-efficient and achieves good balance of performance, latency and cost.
Table 5: Rewriting to balance the quality and latency.
5.3 Performance in Production
As mentioned above, we have developed two pipelines: 1) the detection-only pipeline and 2) the detection with mitigation pipeline. They are integrated into LLM-based products to mitigate the customer’s complaints about hallucination. In practice, we start with the detection-only pipeline and block the hallucinated context from affecting the customers. Gradually, we moved on to the full pipeline of detection and mitigation in the user experience as shown in Figure 4.
Figure 4: User experience of Hallucination Detection & Mitigation Pipeline in Production
Accurately measuring the effectiveness of hallucination detection and mitigation in a production environment is a complex task. We adopt two approaches: we conduct offline human evaluation on some benchmark datasets and we develop a mirror traffic-based flighting system that leverages GPT4 for online quality measurement. This online system applies mirrored traffic to various pipelines and uses GPT4 to assess hallucination rates and the overall quality of rewritten content, and it better monitors the dynamic nature of real-world applications.
The human evaluation result in Table 6 shows that our detection pipelines are very effective. We experimented on key point generation datasets which contain 469 and 161 documents separately. On both datasets, the proposed methods are able to detect hallucinations with precision of above 50% and reduces more than half of the detected hallucinations with precision of above 60%.
Table 6: Offline human evaluation of detection and mitigation.
When applied to a real product scenario, the precision cannot be evaluated directly as we cannot check customer data. As shown by the online result in Table 7, the GPT4-based mitigation can greatly reduce the hallucinated responses and thus the block rate at expense of acceptable increase of latency.
Table 7: Online monitoring and comparisons of the pipelines based on flighting traffic.
6 Challenges, Learnings & Limitations
We have encountered a number of considerable challenges when we integrate this federated hallucination detection and mitigation service hallucination into real products. We have suggested various workable solutions, although some are imperfect, and we outline them with the aim of assisting the community in acknowledging and addressing these challenges collaboratively.
6.1 Measurement of Effectiveness in Production
Accurately measuring the effectiveness of hallucination detection and mitigation and the values bringing to customers in a production environment is very challenging. We have designed a system which applies mirrored traffic to various pipelines and uses GPT4 to assess hallucination rates and the overall quality of rewritten content. However, the GPT4-based measurement has limitation and ensuring the reliability of these measurements and their alignment with human judgment remains an ongoing challenge, necessitating continuous refinement and validation.
6.2 Handling Multilingual and Long Source Documents
We have add several major non-English training data in NLI and SBD models to support the multilingual use case and using segmentation based approach to deal with long source documents. However, managing multilingual even cross-lingual inputs and long source documents adds another layer of complexity. Also it is still an open research problems how to do effective modeling of the long source document.
6.3 Deep Customization Needs for Various Scenarios
Different usages and user circumstances call for tailored adjustments. In the production, we’ve calibrated our ensemble-based detection for greater accuracy with a reduced block rate, while also adjusting our rewriting for decreased latency at the cost of some mitigation power. As we move forward, incorporating additional domain or task-specific datasets may enhance our detection capabilities. Concurrently, researching novel detection techniques or focusing exclusively on detecting certain types or severe cases of hallucinations might be more practical for real-world applications. Furthermore, we should consider investigating more efficient mitigation strategies involving smaller language models.
7 Conclusions
To conclude, we developed a reliable and high-speed production service to detect and mitigate intrinsic or close-world hallucinations in LLMs. Our service combines named entity recognition (NER), natural language inference (NLI), span-based detection (SBD) models, and a decision tree analysis to identify diverse set of hallucinations with high accuracy. We also develop an efficient rewriting mechanism to ensure the quality and fidelity of LLM outputs.
Hallucination is a key problem facing by LLM applications. Our journey from idea to deployment revealed several challenges in delivering high-quality solutions under real-world constraints. We had to balance latency, availability, and quality in production settings.
Our work has broad applicability and potential for future impact, as LLMs are used in diverse sectors that require robust hallucination verification mechanisms. Our service can adapt to different LLM applications, each with different challenges and demands. A key next step is to address extrinsic or open-world hallucinations, which require integrating external knowledge bases and information verification systems.
8 Ethical Considerations
Ethical considerations are paramount in the development and deployment of hallucination mitigation systems. Ensuring transparency in detection and mitigation processes, providing clear explanations for decisions, and safeguarding user data are essential components of our ethical framework. Balancing these ethical imperatives with technical and operational demands is a complex but necessary challenge.
Appendix A Appendix
This is an appendix.